OpenStudy (anonymous):

find the remainder when p(x)=x^3-6x^2+2x-4 when dividedby 1-2x

OpenStudy (anonymous):

i dont remember quite well

OpenStudy (anonymous):

what u remember tell that

OpenStudy (anonymous):

Please reply, you need to take the initiative to reply because its ur work not mine

OpenStudy (anonymous):

or should i tell you what is it

OpenStudy (anonymous):

The remainder theorem is based on synthetic division, which is the process of dividing a polynomial f(x) by a polynomial D(x) and finding the remainder.

OpenStudy (anonymous):

yes!!!

OpenStudy (anonymous):

If p(x) is divided by x-k until a constant remainder R is obtained, then p(k)=R. Proof: p(x) = q(x)(x-k) + R p(k) = q(k)(k-k) + R p(k) = R So -2x + 1 = 0 x = 1/2 p(1/2) = (1/2)^3 - 6(1/2)^2 +2(1/2) - 4 = -35/8 or -4.375

OpenStudy (anonymous):

so how to solve this question b'coz the 1st polynomial of the divisor is bigger than that of the dividend . It's confusing

OpenStudy (anonymous):

that's the remainder -35/8 or -4.375

OpenStudy (anonymous):

can u show it by division method

OpenStudy (anonymous):

it would come the same wait

OpenStudy (anonymous):

would u write the division please?

OpenStudy (anonymous):

wait just 1 minute

OpenStudy (anonymous):

If i do the division , it won't fit in the drawing board @amistre64 any suggestions

OpenStudy (amistre64):

better to walk you thru the division, instead of writing it out for you

OpenStudy (anonymous):

what does it mean?

OpenStudy (amistre64):

first row is the poly with all degrees filled in .... second row is the addons starting with 0 third row is multiplication then its just a repetitive process.

OpenStudy (anonymous):

ms/mr no name is it so long?

OpenStudy (anonymous):

No , it is very simple a intelligent guy like you would be able to do it easily. To represent the division on OS is not possible. So ask your teacher on How to do THE LONG DIVISION OF POLYNOMIALS

OpenStudy (anonymous):

I am not a guy.can I WRITE A PHYSICS QUESTION IN THIS REPLY BOX?

OpenStudy (amistre64):

3 2 1 0, all degrees are present x^3 -6x^2 +2x -4 starting with 0, we add down the columns, just focusing in the coefficients x^3 -6x^2 +2x -4 0 ---------------------- 1/2 | 1 now we multiply the result by the our 1/2 and add it to the next column x^3 -6x^2 +2x -4 0 1/2 ---------------------- 1/2 | 1 -11/2 and repeat till we get to the end

OpenStudy (anonymous):

i cant understand this division it is not simple and of a high level

OpenStudy (amistre64):

its just multiplication and addition .....

OpenStudy (anonymous):

sorry but..........

OpenStudy (anonymous):

See this site and ye will worry no more http://www.wikihow.com/Divide-Polynomials-Using-Synthetic-Division

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

Read everything don't miss out anything please

OpenStudy (anonymous):

If you don't you are always welcome to ask :)

OpenStudy (anonymous):

can I ask a physics question?

OpenStudy (anonymous):

This is for a method of log division:- (Read both methods) http://www.purplemath.com/modules/polydiv2.htm

OpenStudy (anonymous):

Tag me in the physics section , If i can I shall answer

OpenStudy (anonymous):

I am new here so i don't know much about this site.

OpenStudy (anonymous):

are u a teacher no name

OpenStudy (anonymous):

Why ?

OpenStudy (anonymous):

then